1. What is the average Magnetic Flux through one single coil at the initial position of the bar magnet using \bar{B} initial = 2.32 x 10-3 T and \bar{A} = 2.292 x 10-4 m2.

2. What is the total average Magnetic Flux through all 1600 turns in the coil at the initial position of the bar magnet? This is the value for \Phi initial.

3a. What is the average Magnetic Flux through one single coil?

b. What is the total average Magnetic Flux through all 1600 turns in the coil at the final position of the bar magnet? This is the value for \Phi final and it should be x 10-3 of the

\Phiinitial value.
Answer #1

Magnetic flop & BA 2 BA Caso oso $- BA= unren B = 2.32 X1003 T, A=2.292.810-4 m2 $= (2-32 X10-3) X(22.292.X70-4) weber 5.3174